OpenStudy (sriram):

a straight line with negative slope passes through the point (9,4) and cuts the positive co ordinate at P and Q respectively i) min value of OP+OQ (where O is origin) ii)Area of OPQ when OP+OQ is min

OpenStudy (sriram):

options for i) 18 25 36 49

OpenStudy (sriram):

options for ii) 75 225 125 275

OpenStudy (anonymous):

looks like calc to me yes?

OpenStudy (sriram):

we were asked this in co ordinate geo test paper!!

OpenStudy (anonymous):

the points are say (0,y) and (x,0) and you know they have to pass through (9,4) so the slopes must be the same. that way you can write one in terms of the other.

OpenStudy (anonymous):

really? find the max w/o calc? ok lets see

OpenStudy (anonymous):

slope from (9,4) to (x,0) is \[\frac{-4}{x-9}\] and slope from (0,y) to (9,4) is \[\frac{y-4}{-9}\] and these have to be equal so i guess \[\frac{-4}{x-9}=\frac{y-4}{-9}\] maybe we can solve for one in terms of the other

OpenStudy (sriram):

i tried to assume a slope m y-4=m(x-9) x intercept=9-4/m and y intercept=4-9m

OpenStudy (sriram):

OP+OQ=9-4/m+4-9m =9m-4+4m-9m^2 ---------------- m m is negative

OpenStudy (anonymous):

oh this is probably better than what i was trying \[x=9-\frac{4}{m}\] \[y =4-9m\]

OpenStudy (sriram):

-9m^2+13m-4 ------------- needs to be minimum m do u think differentiating it would be the best way possible?

OpenStudy (anonymous):

wait i got lost somewhere there. when you add you get \[9-\frac{4}{m}+4-9m\] yes?

OpenStudy (anonymous):

oh ok i see you put it all over m \[\frac{-9m^2+13m+4}{m}\]

OpenStudy (anonymous):

ok and we want the minimum value out of this. that will give the slope. looks like you are doing better than me.

OpenStudy (anonymous):

except i think it should be +4 yes, not -4 or did i mess up?

OpenStudy (anonymous):

oh no it is -4

OpenStudy (anonymous):

Can't you just look at this geometrically? You have a line pivoting about (9,4) through pi/2

OpenStudy (sriram):

u mean the angle wid x axis as pi/2?

OpenStudy (anonymous):

ok derivative is \[\frac{4-9x^2}{x^2}\] set equal zero get \[9x^2=4\] \[x=-\frac{2}{3}\]

OpenStudy (anonymous):

since x is negative.

OpenStudy (anonymous):

From min to max, the slope (Tan) varies from y=9+, x is 4+

OpenStudy (anonymous):

put in \[-4=-\frac{2}{3}(x-9)\] gives \[x=15\]

OpenStudy (anonymous):

so x intercept is (15,0) and then y = 10

OpenStudy (anonymous):

wow you had it. better than my method.

OpenStudy (sriram):

can u jst show the working of how u got the dervatve coz i m getting a cubic differentiating it

OpenStudy (anonymous):

sure but why don't we do it the easy way first?

OpenStudy (anonymous):

we actually started with \[-9m+16-\frac{4}{m}\] before we put over one denominator. so just take the derivative of this thing

OpenStudy (anonymous):

get \[-9+\frac{4}{m^2}\] right away

OpenStudy (sriram):

oh and i wasted 15 minutes in the paper breaking my head over the cubic!!!!!

OpenStudy (anonymous):

or \[\frac{4-9m^2}{m^2}\] which is what you will get if you use the quotient rule. want to do it that way too?

OpenStudy (anonymous):

yeah but it will work out the same with more algebra for sure. can we skip that or would you like to do it as well? up to you

OpenStudy (sriram):

i would love to see the algebric way of solving it

OpenStudy (anonymous):

well first of all you do not get a cubic. we have \[f(m)=-9m^2+13m-4\] \[f'(m)=-18m+13\] \[g(m)=m\] \[g'(m)=1\] and use \[\frac{f}{g}'=\frac{gf'-f'g}{g^2}\]

OpenStudy (anonymous):

gives \[\frac{(-18m+13)-1(-9m^2+13m-4)}{m^2}\]

OpenStudy (anonymous):

ok that was a mistake

OpenStudy (anonymous):

\[\frac{m(-18m+13)-1(-9m^2+13m-4)}{m^2}\] is better

OpenStudy (anonymous):

numerator is \[-18m^2+13m+9m^2-13m+4=-9m^2+4\]

OpenStudy (sriram):

yep u got it

OpenStudy (anonymous):

and that gets it

OpenStudy (sriram):

thanks a lot

OpenStudy (anonymous):

someone help me lol please :) you guys are really smart !

OpenStudy (anonymous):

but you did all the work. the rest was mechanics. solution is x = 15, y = 10 and done

OpenStudy (anonymous):

@tabby choice c

OpenStudy (sriram):

to tabbeyx33 it is the 3rd option

OpenStudy (anonymous):

thanksthanks <3333 lov yaa :)

OpenStudy (anonymous):

so let me scroll up and see what we got, because if it is not there i am going to be bummed.

OpenStudy (anonymous):

yup 25 for the sum and 75 for the product. good work

OpenStudy (anonymous):

so last one guys :) its another type of question as the previous i just asked

OpenStudy (anonymous):

b

OpenStudy (anonymous):

thank you soo much <3 i passed :) wow your like super awesome :)

OpenStudy (sriram):

b it is

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Still puzzling over the geometry of this. In particular, why is it that the slope for the minimum squared = 4/9? Not a happy coincidence, surely.

OpenStudy (anonymous):

If we let y be the difference beween the y-intercept and 9 and x the difference between the x-intercept and 4, we have tangent (slope) = y/4 = 9/x We want Min(x + y +13) = Min(x + 36/x +13) Diff = 1 - 36/x^2 -> x = 6 (-6 not) which is evidently the same if you use y instead. So you have the same 10 by 15 axes reversed.

OpenStudy (anonymous):

no i was fairly sure it was not a coincidence that the slope turned out to be 2/3